OpenStudy (anonymous):

8e^3x+9=13. Round to the nearest thousandth

OpenStudy (kirbykirby):

\[ 8e^{3x}+9=13\\ 8e^{3x}=4\\ e^{3x}=\frac{4}{8}=\frac{1}{2}\\ \ln(e^{3x})=\ln\left( \frac{1}{2}\right)=-\ln 2\\ 3x(\ln e)=-\ln 2\\ 3x=-\ln 2\\ \, \\ x = \frac{-\ln 2}{3}\approx-0.231\]